Let M = f(t) be the average number of meals per person purchased at restaurants annually in
the year that is t years since 1900. An equation of f is f(t) = 2.22t – 71.01.
a. Find an equation for f^–1.
b. Find f(125)
c. Find f^–1 (125)
Please solve a,b, and c for the question above.
Please show me step by step to get the answer. Thanks
I really DO NOT know how to solve this problem. Please HELP ME!! Thanks

Answer by stanbon(75887)

Let M = f(t) be the average number of meals per person purchased at restaurants annually in
the year that is t years since 1900.
An equation of f is f(t) = 2.22t – 71.01.
y = 2.22t-71.01
-------
a. Find an equation for f^–1.
1st: Interchange t and y to get:
t = 2.22y-71.01
2nd: Solve for "y":
f^-1(t) = (t+71.01)/2.22
-----
b. Find f(125) = 2.22*125 - 71.01 = 206.49
----------
c. Find f^–1 (125) = (125+71.01)/2.22 = 88.2928
=============================
Cheers,
Stan H.